# Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning

### v2

@@ 1 Introduction
This naturally suggests a sequential decision problem: choose the next propagation to maximize long-horizon gains in pruning and solution quality under compute budgets. Reinforcement learning has repeatedly shown that learned policies can outperform handcrafted heuristics in domains where actions have delayed, global consequences (mnih2015human; silver2016alphago; silver2017alphago; lillicrap2015continuous; watkins1992q; williams1992simple). Here, each propagation step becomes an action conditioned on the current solver state, enabling the policy to trade off immediate cost against future reductions in search and error (rumelhart1986learning; lecun1998gradient; hochreiter1997long).
We introduce MetaJuLS, a meta-reinforcement learning approach for transferable adaptive constraint propagation that scales to modern LLM inference. We represent the inference state as a constraint–variable graph and parameterize the scheduler with a Graph Attention Network (velickovic2018graph). To generalize across languages and constraint types, we meta-train the policy with Model-Agnostic Meta-Learning (finn2017model), enabling few-step adaptation (5–10 gradient steps; 5–15 seconds) rather than hours of task-specific retraining. We use “zero-shot” strictly for transfer with no gradient steps; our setting is rapid few-shot adaptation.
@@ 2.3 Related Work
Our work builds on three research threads: learned heuristics for constraint programming (gasse2019exact; cappart2021combinatorial), graph neural networks for structured optimization (velickovic2018graph; kipf2017semi), and reinforcement learning for combinatorial problems (sutton2018reinforcement; schulman2017proximal). Unlike prior work on search heuristics (gasse2019exact) or reactive scheduling (marques2012activity), MetaJuLS learns predictive propagation policies via meta-RL that generalize across languages and constraint types. A detailed survey appears in Appendix A.
@@ 3.1 Problem Formulation
Formally, the MDP is defined by the tuple , where represents the set of possible solver states, is the action space of constraint propagators, defines the transition dynamics, is the reward function, and is the discount factor (sutton2018reinforcement). The state space encodes the current domains of all variables, the constraint graph structure, and metadata about recent propagation activity (bessiere2006constraint). Each state is represented as a graph , where vertices correspond to variables and constraints, and edges represent variable-constraint relationships.
The action space consists of all constraint propagators that are currently “dirty,” that is, constraints whose associated variables have experienced domain changes since their last propagation (bessiere2006constraint). At each time step , the agent selects an action , where is the subset of dirty constraints at time . The selected propagator is then executed, potentially reducing variable domains and marking additional constraints as dirty for future propagation (regin1994filtering; lecoutre2009str).
@@
The reward function balances the domain reduction achieved by propagating constraint against the computational cost of the propagation operation (sutton2018reinforcement). Specifically, we define the reward as:
@@
where is the set of variables involved in constraint , is the current domain size of variable , and is a constraint-type-specific constant reflecting the complexity of the filtering algorithm. The constants reflect empirical operation counts: grammar rule constraints require checking two child domains (binary lookup), while lexical constraints require a single vocabulary check. These values are robust to perturbation; Appendix C.1 reports 2% performance variation for . For grammar rule constraints in parsing, we set (binary checks), while for lexical constraints (simpler lookups). For CP constraints, for binary constraints and for alldiff where is arity. This deterministic proxy avoids hardware-dependent wall-clock time measurements while accurately reflecting computational cost (schulman2017proximal). To validate the cost proxy, we measure correlation between proxy cost and wall-clock time across constraint types and sentence lengths. On PTB parsing, we find Pearson correlation (, ) between proxy cost and measured microseconds per propagator, confirming that optimizing the proxy improves real wall-clock time. The reward structure encourages the agent to discover policies that maximize search space reduction while minimizing solver overhead (debruyne1997optimal; gent2006watched).
The transition dynamics are deterministic given the constraint propagator semantics, but the state space is extremely large and the effects of propagation can cascade through the constraint graph in complex ways (apt2003principles; marriott1998programming). The discount factor is set close to 1 (typically 0.99) to encourage long-term planning, as the benefits of effective propagation scheduling may only become apparent after many propagation steps (sutton2018reinforcement; schulman2017proximal).
@@ 3.2 Architecture: Graph Attention Network
To encode the solver state and enable effective action selection, we employ a Graph Attention Network (GAT) architecture (velickovic2018graph) that processes the constraint-variable graph structure. The GAT allows the agent to reason about how domain reductions propagate through the constraint network, identifying bottleneck constraints and prioritizing propagators that will have the greatest downstream impact.
The input to the GAT is a graph where each node represents either a variable or a constraint (velickovic2018graph; kipf2017semi; gehring2017convolutional). For variable nodes, we include features such as the current domain size, the number of constraints involving the variable, and recent domain change history (bessiere2006constraint). For constraint nodes, we include features such as the constraint type, the number of variables involved, the current violation magnitude, and activity scores based on recent propagation behavior (marques2012activity; moskewicz2001chaff).
@@
where and are the feature vectors of nodes and (velickovic2018graph), is a learned weight matrix, is a learned attention vector (vaswani2017attention), is the set of neighbors of node , and denotes concatenation. The node representations are then updated as:
where is a nonlinear activation function (lecun1998gradient; hochreiter1997long).
After processing through multiple GAT layers, we obtain enriched node representations that capture both local constraint-variable relationships and global graph structure (hamilton2017inductive; perozzi2014deepwalk). For action selection, we apply a final attention layer that computes scores for each constraint node, indicating the expected value of propagating that constraint (vaswani2017attention). The constraint with the highest score is selected as the next action, or we sample from a softmax distribution over scores for exploration during training (schulman2017proximal).
@@ 3.3 Meta-Learning for Universal Generalization
To enable rapid adaptation across languages and tasks, we train MetaJuLS using Model-Agnostic Meta-Learning (MAML) (finn2017model). Rather than training a single policy for one task, we meta-train over a distribution of tasks (e.g., different languages, grammars, or constraint types). The meta-objective learns initial parameters that can quickly adapt to new tasks with only 5–10 gradient steps, requiring seconds rather than hours of training.
@@ 3.4 Training Procedure
We combine meta-learning with Proximal Policy Optimization (PPO) (schulman2017proximal) for stable policy updates. The inner-loop loss uses the standard PPO clipped objective. During meta-training, we generate episodes across diverse tasks: parsing in 10 languages (Universal Dependencies), constrained decoding with different schemas, and CP problems from multiple domains. For each task, we collect trajectories and compute advantages using Generalized Advantage Estimation (GAE) (schulman2017proximal). The meta-training procedure alternates between inner-loop adaptation (fast task-specific learning) and outer-loop meta-updates (learning to learn), enabling the policy to generalize to unseen tasks.
The training objective combines the policy gradient term with a value function loss and an entropy bonus to encourage exploration (schulman2017proximal):
where is the importance sampling ratio, is the estimated advantage (sutton2018reinforcement), is the value function loss, is the policy entropy, and and are hyperparameters (schulman2017proximal).
We train for multiple epochs over collected trajectories, updating the policy network using the Adam optimizer (kingma2014adam) with a learning rate of (lecun2015deep). To improve sample efficiency while preserving the on-policy nature of PPO, we perform several gradient update epochs on each freshly collected batch of trajectories before discarding it and collecting new data (mnih2015human). Training typically converges after 50–100 epochs, depending on the complexity of the problem domains (goodfellow2016deep).
@@ 4.1 Constituency Parsing: Penn Treebank
We evaluate MetaJuLS on constituency parsing using the Penn Treebank (PTB) Wall Street Journal corpus (marcus1993building), following the standard split: sections 2–21 for training, section 22 for development, and section 23 for testing. We implement CKY-style chart parsing as described in Section 2, where variables correspond to spans, domains contain candidate nonterminal labels, and constraints encode grammar rules. All comparisons use the same base parser, beam size, and batching; only the constraint scheduling policy differs.
We compare against Berkeley Neural Parser (kitaev2018constituency), GPU-parallelized CKY (our reimplementation of klein2003accurate using batched CUDA kernels for span-parallel chart filling), A* search with neural heuristics (stern2017minimum), activity-based scheduling, VSIDS-style conflict-driven scheduling (prioritizing constraints that recently caused pruning), cost-normalized greedy (maximizing using our deterministic cost proxy), and supervised ranking (learning-to-rank constraints from state features, trained to imitate oracle best-first on development set). The meta-learned MetaJuLS policy is trained on PTB sections 2–21. Table 1 reports F1 scores and inference time. MetaJuLS achieves 1.6–1.9 speedups over GPU-parallelized CKY while maintaining within 0.2% F1 of Berkeley Neural Parser. MetaJuLS speedups increase with sentence length, achieving 1.91 on sentences tokens vs. 1.2 on short sentences (full scaling analysis in Appendix B.2).
@@ 4.2 Dependency Parsing: Universal Dependencies
We evaluate MetaJuLS on dependency parsing using Universal Dependencies (UD) v2.11 (zeman2022universal) across 10 diverse languages: English, Spanish, French, German, Chinese, Arabic, Finnish, Japanese, Russian, and Hindi. We implement a transition-based dependency parser where variables correspond to parser states, domains encode possible transitions, and constraints enforce arc consistency and well-formedness. All UD comparisons use the same dependency parser architecture, beam size, and batching; only the constraint scheduling policy changes, ensuring apples-to-apples comparison.
We compare against state-of-the-art dependency parsers (UDPipe (straka2016udpipe), Stanza (qi2020stanza)), activity-based scheduling, VSIDS-style conflict-driven scheduling, cost-normalized greedy, and MetaJuLS. The meta-learned MetaJuLS policy is trained on English UD data, then rapidly adapted to other languages with 5–10 gradient steps (requiring 5–15 seconds). Table 2 reports LAS (Labeled Attachment Score) and inference time. MetaJuLS achieves 1.5–1.8 speedups over activity-based scheduling while maintaining competitive LAS scores, as shown in Figure 2.
@@ 4.3 LLM Constrained Decoding
We evaluate MetaJuLS on large language model inference where outputs must satisfy formal constraints. We use Llama-2-7B (touvron2023llama) with vLLM (kwon2023efficient) for efficient serving, integrating MetaJuLS into the decoding loop to enforce JSON schemas and logical constraints.
@@
We evaluate on LogicBench (chen2024logicbench), which requires generating formal logical expressions that satisfy given premises. Variables correspond to token positions, domains contain logical operators and predicates, and constraints enforce well-formedness and logical consistency. To test whether MetaJuLS and speculative decoding are complementary, we perform a factorial experiment comparing: (1) baseline (neither method), (2) speculative decoding only, (3) MetaJuLS scheduling only, and (4) both methods combined. We compare against NeuroLogic Decoding (lu2021neurologic), Outlines (willard2023outlines), and LMQL (beurer2023lmql). MetaJuLS achieves 1.8 speedup over speculative decoding alone while maintaining 98.2% constraint satisfaction rate (vs. 94.1% for speculative decoding). When combined, speculative decoding + MetaJuLS achieves an additional 15% speedup over MetaJuLS alone, demonstrating orthogonality.
On GSM8K (cobbe2021training) with JSON schema constraints requiring structured answer formats, MetaJuLS accelerates generation by 1.6 compared to speculative decoding while improving schema compliance from 91.3% to 96.8%. Table 3 summarizes these results.
@@ 6 Discussion and Conclusion
MetaJuLS shows that reinforcement learning can learn adaptive constraint-propagation schedules that outperform handcrafted and reactive heuristics by treating scheduling as a sequential decision problem with delayed, global effects (silver2016alphago; silver2017alphago; schulman2017proximal; sutton2018reinforcement; bessiere2006constraint; debruyne1997optimal; shihab2025detecting). Policy analysis indicates that the learned scheduler both rediscovers effective classical principles in simple regimes (e.g., Domain-over-Weight) and discovers non-trivial strategies in harder structured settings (e.g., prioritizing violation potential), supporting the claim that learning can capture and extend beyond standard heuristic design (apt2003principles; marriott1998programming; regin1994filtering; gent2006watched).
@@ Appendix A Extended Related Work
The intersection of machine learning and constraint programming has received increasing attention in recent years, with researchers exploring learned heuristics for variable ordering, value selection, and constraint propagation (gasse2019exact; cappart2021combinatorial; nair2020solving). Early work in this area focused on supervised learning approaches, training classifiers to predict which constraints should be propagated next based on handcrafted features (wallace1996practical; benhamou2004interval; andrew2007scalable). While these methods showed promise, they suffered from the fundamental limitation that optimal propagation decisions depend on the entire search trajectory, not just local features (Ando2005; rasooli-tetrault-2015).
@@
Activity-based heuristics, inspired by the VSIDS (Variable State Independent Decaying Sum) strategy from SAT solving (moskewicz2001chaff; marques2012activity), represent a significant advance over static scheduling (regin1994filtering; lecoutre2009str). These methods maintain activity scores for constraints based on their historical contribution to search progress, dynamically prioritizing constraints that have recently caused domain reductions or failures (audemard2013glucose; debruyne1997optimal). However, activity-based heuristics remain reactive rather than predictive, adjusting priorities only after observing constraint behavior rather than anticipating which constraints will be most effective in the current context.
@@
Graph neural networks have emerged as a powerful tool for learning representations of structured data, with applications ranging from molecular property prediction to social network analysis (kipf2017semi; hamilton2017inductive; perozzi2014deepwalk; vaswani2017attention). Most directly related is the line of work applying graph neural networks to combinatorial optimization. gasse2019exact learn branching policies for mixed-integer programming using GNNs over the bipartite variable-constraint graph, demonstrating that learned policies can match or exceed SCIP’s default heuristics. cappart2021combinatorial survey the broader landscape of machine learning for constraint programming, identifying propagation scheduling as an underexplored direction. nair2020solving apply GNNs to SAT solving, learning clause selection policies. In constraint programming, recent work has explored using graph neural networks to learn variable ordering heuristics (velickovic2018graph), demonstrating that learned policies can outperform classical heuristics on specific problem classes (gent2006watched). However, these approaches have primarily focused on search heuristics rather than propagation scheduling, and have not addressed the challenge of generalizing across diverse problem domains. Our work extends this paradigm to NLP inference, demonstrating that policies learned on linguistic constraints transfer to classical CP benchmarks and vice versa—a bidirectional generalization not previously established.
@@
Reinforcement learning has shown remarkable success in combinatorial optimization, with applications to vehicle routing, scheduling, and resource allocation (sutton2018reinforcement; schulman2017proximal; freund1997decision). The key insight from this line of work is that many optimization problems can be naturally formulated as sequential decision processes, where an agent learns to construct solutions incrementally through trial and error. In constraint programming, this perspective suggests that propagation scheduling might benefit from a similar treatment, with the agent learning to construct effective propagation sequences through interaction with the solver.
@@
Our work builds on these foundations by combining graph neural networks with reinforcement learning to learn adaptive propagation policies (velickovic2018graph; schulman2017proximal). Unlike previous approaches that rely on handcrafted features or reactive heuristics (wallace1996practical; marques2012activity), MetaJuLS learns to extract relevant information from the constraint-variable graph structure itself (kipf2017semi; hamilton2017inductive), enabling the discovery of novel scheduling strategies that adapt to both problem structure and solver state (bessiere2006constraint).

### v3

@@ 1 Introduction
This naturally suggests a sequential decision problem: choose the next propagation to maximize long-horizon gains in pruning and solution quality under compute budgets. Reinforcement learning has repeatedly shown that learned policies can outperform handcrafted heuristics in domains where actions have delayed, global consequences (Mnih et al., 2015; Silver et al., 2016, 2017; Lillicrap et al., 2015; Watkins and Dayan, 1992; Williams, 1992). Here, each propagation step becomes an action conditioned on the current solver state, enabling the policy to trade off immediate cost against future reductions in search and error (Rumelhart et al., 1986; LeCun et al., 1998; Hochreiter and Schmidhuber, 1997).
We introduce MetaJuLS, a meta-reinforcement learning approach for transferable adaptive constraint propagation that scales to modern LLM inference. We represent the inference state as a constraint–variable graph and parameterize the scheduler with a Graph Attention Network (Veličković et al., 2018). To generalize across languages and constraint types, we meta-train the policy with Model-Agnostic Meta-Learning (Finn et al., 2017), enabling few-step adaptation (5–10 gradient steps; 5–15 seconds) rather than hours of task-specific retraining. We use “zero-shot” strictly for transfer with no gradient steps; our setting is rapid few-shot adaptation.
@@ 2.3 Related Work
Our work builds on three research threads: learned heuristics for constraint programming (Gasse et al., 2019; Cappart et al., 2021), graph neural networks for structured optimization (Veličković et al., 2018; Kipf and Welling, 2017), and reinforcement learning for combinatorial problems (Sutton and Barto, 2018; Schulman et al., 2017). Unlike prior work on search heuristics (Gasse et al., 2019) or reactive scheduling (Marques-Silva et al., 2012), MetaJuLS learns predictive propagation policies via meta-RL that generalize across languages and constraint types. A detailed survey appears in Appendix A.
@@ 3.1 Problem Formulation
Formally, the MDP is defined by the tuple , where represents the set of possible solver states, is the action space of constraint propagators, defines the transition dynamics, is the reward function, and is the discount factor (Sutton and Barto, 2018). The state space encodes the current domains of all variables, the constraint graph structure, and metadata about recent propagation activity (Bessière, 2006). Each state is represented as a graph , where vertices correspond to variables and constraints, and edges represent variable-constraint relationships.
The action space consists of all constraint propagators that are currently “dirty,” that is, constraints whose associated variables have experienced domain changes since their last propagation (Bessière, 2006). At each time step , the agent selects an action , where is the subset of dirty constraints at time . The selected propagator is then executed, potentially reducing variable domains and marking additional constraints as dirty for future propagation (Régin, 1994; Lecoutre, 2009).
@@
The reward function balances the domain reduction achieved by propagating constraint against the computational cost of the propagation operation (Sutton and Barto, 2018). Specifically, we define the reward as:
@@
where is the set of variables involved in constraint , is the current domain size of variable , and is a constraint-type-specific constant reflecting the complexity of the filtering algorithm. The constants reflect empirical operation counts: grammar rule constraints require checking two child domains (binary lookup), while lexical constraints require a single vocabulary check. These values are robust to perturbation; Appendix C.1 reports 2% performance variation for . For grammar rule constraints in parsing, we set (binary checks), while for lexical constraints (simpler lookups). For CP constraints, for binary constraints and for alldiff where is arity. This deterministic proxy avoids hardware-dependent wall-clock time measurements while accurately reflecting computational cost (Schulman et al., 2017). To validate the cost proxy, we measure correlation between proxy cost and wall-clock time across constraint types and sentence lengths. On PTB parsing, we find Pearson correlation (, ) between proxy cost and measured microseconds per propagator, confirming that optimizing the proxy improves real wall-clock time. The reward structure encourages the agent to discover policies that maximize search space reduction while minimizing solver overhead (Debruyne and Bessière, 1997; Gent et al., 2006).
The transition dynamics are deterministic given the constraint propagator semantics, but the state space is extremely large and the effects of propagation can cascade through the constraint graph in complex ways (Apt, 2003; Marriott and Stuckey, 1998). The discount factor is set close to 1 (typically 0.99) to encourage long-term planning, as the benefits of effective propagation scheduling may only become apparent after many propagation steps (Sutton and Barto, 2018; Schulman et al., 2017).
@@ 3.2 Architecture: Graph Attention Network
To encode the solver state and enable effective action selection, we employ a Graph Attention Network (GAT) architecture (Veličković et al., 2018) that processes the constraint-variable graph structure. The GAT allows the agent to reason about how domain reductions propagate through the constraint network, identifying bottleneck constraints and prioritizing propagators that will have the greatest downstream impact.
The input to the GAT is a graph where each node represents either a variable or a constraint (Veličković et al., 2018; Kipf and Welling, 2017; Gehring et al., 2017). For variable nodes, we include features such as the current domain size, the number of constraints involving the variable, and recent domain change history (Bessière, 2006). For constraint nodes, we include features such as the constraint type, the number of variables involved, the current violation magnitude, and activity scores based on recent propagation behavior (Marques-Silva et al., 2012; Moskewicz et al., 2001).
@@
where and are the feature vectors of nodes and (Veličković et al., 2018), is a learned weight matrix, is a learned attention vector (Vaswani et al., 2017), is the set of neighbors of node , and denotes concatenation. The node representations are then updated as:
where is a nonlinear activation function (LeCun et al., 1998; Hochreiter and Schmidhuber, 1997).
After processing through multiple GAT layers, we obtain enriched node representations that capture both local constraint-variable relationships and global graph structure (Hamilton et al., 2017; Perozzi et al., 2014). For action selection, we apply a final attention layer that computes scores for each constraint node, indicating the expected value of propagating that constraint (Vaswani et al., 2017). The constraint with the highest score is selected as the next action, or we sample from a softmax distribution over scores for exploration during training (Schulman et al., 2017).
@@ 3.3 Meta-Learning for Universal Generalization
To enable rapid adaptation across languages and tasks, we train MetaJuLS using Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017). Rather than training a single policy for one task, we meta-train over a distribution of tasks (e.g., different languages, grammars, or constraint types). The meta-objective learns initial parameters that can quickly adapt to new tasks with only 5–10 gradient steps, requiring seconds rather than hours of training.
@@ 3.4 Training Procedure
We combine meta-learning with Proximal Policy Optimization (PPO) (Schulman et al., 2017) for stable policy updates. The inner-loop loss uses the standard PPO clipped objective. During meta-training, we generate episodes across diverse tasks: parsing in 10 languages (Universal Dependencies), constrained decoding with different schemas, and CP problems from multiple domains. For each task, we collect trajectories and compute advantages using Generalized Advantage Estimation (GAE) (Schulman et al., 2017). The meta-training procedure alternates between inner-loop adaptation (fast task-specific learning) and outer-loop meta-updates (learning to learn), enabling the policy to generalize to unseen tasks.
The training objective combines the policy gradient term with a value function loss and an entropy bonus to encourage exploration (Schulman et al., 2017):
where is the importance sampling ratio, is the estimated advantage (Sutton and Barto, 2018), is the value function loss, is the policy entropy, and and are hyperparameters (Schulman et al., 2017).
We train for multiple epochs over collected trajectories, updating the policy network using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of (LeCun et al., 2015). To improve sample efficiency while preserving the on-policy nature of PPO, we perform several gradient update epochs on each freshly collected batch of trajectories before discarding it and collecting new data (Mnih et al., 2015). Training typically converges after 50–100 epochs, depending on the complexity of the problem domains (Goodfellow et al., 2016).
@@ 4.1 Constituency Parsing: Penn Treebank
We evaluate MetaJuLS on constituency parsing using the Penn Treebank (PTB) Wall Street Journal corpus (Marcus et al., 1993), following the standard split: sections 2–21 for training, section 22 for development, and section 23 for testing. We implement CKY-style chart parsing as described in Section 2, where variables correspond to spans, domains contain candidate nonterminal labels, and constraints encode grammar rules. All comparisons use the same base parser, beam size, and batching; only the constraint scheduling policy differs.
We compare against Berkeley Neural Parser (Kitaev and Klein, 2018), GPU-parallelized CKY (our reimplementation of Klein and Manning (2003) using batched CUDA kernels for span-parallel chart filling), A* search with neural heuristics (Stern et al., 2017), activity-based scheduling, VSIDS-style conflict-driven scheduling (prioritizing constraints that recently caused pruning), cost-normalized greedy (maximizing using our deterministic cost proxy), and supervised ranking (learning-to-rank constraints from state features, trained to imitate oracle best-first on development set). The meta-learned MetaJuLS policy is trained on PTB sections 2–21. Table 1 reports F1 scores and inference time. MetaJuLS achieves 1.6–1.9 speedups over GPU-parallelized CKY while maintaining within 0.2% F1 of Berkeley Neural Parser. MetaJuLS speedups increase with sentence length, achieving 1.91 on sentences tokens vs. 1.2 on short sentences (full scaling analysis in Appendix B.2).
@@ 4.2 Dependency Parsing: Universal Dependencies
We evaluate MetaJuLS on dependency parsing using Universal Dependencies (UD) v2.11 (Nivre et al., 2020) across 10 diverse languages: English, Spanish, French, German, Chinese, Arabic, Finnish, Japanese, Russian, and Hindi. We implement a transition-based dependency parser where variables correspond to parser states, domains encode possible transitions, and constraints enforce arc consistency and well-formedness. All UD comparisons use the same dependency parser architecture, beam size, and batching; only the constraint scheduling policy changes, ensuring apples-to-apples comparison.
We compare against state-of-the-art dependency parsers (UDPipe (Straka and Strakova, 2016), Stanza (Qi et al., 2020)), activity-based scheduling, VSIDS-style conflict-driven scheduling, cost-normalized greedy, and MetaJuLS. The meta-learned MetaJuLS policy is trained on English UD data, then rapidly adapted to other languages with 5–10 gradient steps (requiring 5–15 seconds). Table 2 reports LAS (Labeled Attachment Score) and inference time. MetaJuLS achieves 1.5–1.8 speedups over activity-based scheduling while maintaining competitive LAS scores, as shown in Figure 2.
@@ 4.3 LLM Constrained Decoding
We evaluate MetaJuLS on large language model inference where outputs must satisfy formal constraints. We use Llama-2-7B (Touvron et al., 2023) with vLLM (Kwon et al., 2023) for efficient serving, integrating MetaJuLS into the decoding loop to enforce JSON schemas and logical constraints.
@@
We evaluate on LogicBench (Chen et al., 2024), which requires generating formal logical expressions that satisfy given premises. Variables correspond to token positions, domains contain logical operators and predicates, and constraints enforce well-formedness and logical consistency. To test whether MetaJuLS and speculative decoding are complementary, we perform a factorial experiment comparing: (1) baseline (neither method), (2) speculative decoding only, (3) MetaJuLS scheduling only, and (4) both methods combined. We compare against NeuroLogic Decoding (Lu et al., 2021), Outlines (Willard and Louf, 2023), and LMQL (Beurer-Kellner et al., 2023). MetaJuLS achieves 1.8 speedup over speculative decoding alone while maintaining 98.2% constraint satisfaction rate (vs. 94.1% for speculative decoding). When combined, speculative decoding + MetaJuLS achieves an additional 15% speedup over MetaJuLS alone, demonstrating orthogonality.
On GSM8K (Cobbe et al., 2021) with JSON schema constraints requiring structured answer formats, MetaJuLS accelerates generation by 1.6 compared to speculative decoding while improving schema compliance from 91.3% to 96.8%. Table 3 summarizes these results.
@@ 6 Discussion and Conclusion
MetaJuLS shows that reinforcement learning can learn adaptive constraint-propagation schedules that outperform handcrafted and reactive heuristics by treating scheduling as a sequential decision problem with delayed, global effects (Silver et al., 2016, 2017; Schulman et al., 2017; Sutton and Barto, 2018; Bessière, 2006; Debruyne and Bessière, 1997; Shihab et al., 2026). Policy analysis indicates that the learned scheduler both rediscovers effective classical principles in simple regimes (e.g., Domain-over-Weight) and discovers non-trivial strategies in harder structured settings (e.g., prioritizing violation potential), supporting the claim that learning can capture and extend beyond standard heuristic design (Apt, 2003; Marriott and Stuckey, 1998; Régin, 1994; Gent et al., 2006).
@@ 7 Limitations
Acknowledgments
We thank the anonymous reviewers for their valuable feedback and suggestions. This work was supported by [funding information]. We also acknowledge the developers of the MiniZinc and XCSP benchmark suites for providing standardized evaluation datasets (Nethercote et al., 2011; Lecoutre et al., 2011).
@@ Appendix A Extended Related Work
The intersection of machine learning and constraint programming has received increasing attention in recent years, with researchers exploring learned heuristics for variable ordering, value selection, and constraint propagation (Gasse et al., 2019; Cappart et al., 2021; Nair et al., 2020). Early work in this area focused on supervised learning approaches, training classifiers to predict which constraints should be propagated next based on handcrafted features (Wallace, 1996; Benhamou et al., 2004; Andrew and Gao, 2007). While these methods showed promise, they suffered from the fundamental limitation that optimal propagation decisions depend on the entire search trajectory, not just local features (Ando and Zhang, 2005; Rasooli and Tetreault, 2015).
@@
Activity-based heuristics, inspired by the VSIDS (Variable State Independent Decaying Sum) strategy from SAT solving (Moskewicz et al., 2001; Marques-Silva et al., 2012), represent a significant advance over static scheduling (Régin, 1994; Lecoutre, 2009). These methods maintain activity scores for constraints based on their historical contribution to search progress, dynamically prioritizing constraints that have recently caused domain reductions or failures (Audemard and Simon, 2013; Debruyne and Bessière, 1997). However, activity-based heuristics remain reactive rather than predictive, adjusting priorities only after observing constraint behavior rather than anticipating which constraints will be most effective in the current context.
@@
Graph neural networks have emerged as a powerful tool for learning representations of structured data, with applications ranging from molecular property prediction to social network analysis (Kipf and Welling, 2017; Hamilton et al., 2017; Perozzi et al., 2014; Vaswani et al., 2017). Most directly related is the line of work applying graph neural networks to combinatorial optimization. Gasse et al. (2019) learn branching policies for mixed-integer programming using GNNs over the bipartite variable-constraint graph, demonstrating that learned policies can match or exceed SCIP’s default heuristics. Cappart et al. (2021) survey the broader landscape of machine learning for constraint programming, identifying propagation scheduling as an underexplored direction. Nair et al. (2020) apply GNNs to SAT solving, learning clause selection policies. In constraint programming, recent work has explored using graph neural networks to learn variable ordering heuristics (Veličković et al., 2018), demonstrating that learned policies can outperform classical heuristics on specific problem classes (Gent et al., 2006). However, these approaches have primarily focused on search heuristics rather than propagation scheduling, and have not addressed the challenge of generalizing across diverse problem domains. Our work extends this paradigm to NLP inference, demonstrating that policies learned on linguistic constraints transfer to classical CP benchmarks and vice versa—a bidirectional generalization not previously established.
@@
Reinforcement learning has shown remarkable success in combinatorial optimization, with applications to vehicle routing, scheduling, and resource allocation (Sutton and Barto, 2018; Schulman et al., 2017; Freund and Schapire, 1997). The key insight from this line of work is that many optimization problems can be naturally formulated as sequential decision processes, where an agent learns to construct solutions incrementally through trial and error. In constraint programming, this perspective suggests that propagation scheduling might benefit from a similar treatment, with the agent learning to construct effective propagation sequences through interaction with the solver.
@@
Our work builds on these foundations by combining graph neural networks with reinforcement learning to learn adaptive propagation policies (Veličković et al., 2018; Schulman et al., 2017). Unlike previous approaches that rely on handcrafted features or reactive heuristics (Wallace, 1996; Marques-Silva et al., 2012), MetaJuLS learns to extract relevant information from the constraint-variable graph structure itself (Kipf and Welling, 2017; Hamilton et al., 2017), enabling the discovery of novel scheduling strategies that adapt to both problem structure and solver state (Bessière, 2006).